A car is travelling at constant speed of 60m/s on a highway. Another car needed to refuel so starting from rest, it leaves the area with an acceleration of 7m/2. It takes 3s for this car to enter the highway. This car keeps its acceleration constant on the highway. How long does it take for the entering car to catch the one that is travelling at 60m/s?  
Answer #1

Distance travelled by the car at Constant speed

X1=VT =60T

Speed of the second car after t=3 s ,from

V=Vo+at

V=0+7*3 =21 m/s-----------------1

The Distance traveled by second car in "T" seconds

X2=Vot+(1/2)at2

X2 = 21t+(1/2)(7)t2

X2 = 21T+3.5T2-----------------2

Therefore

X1=X2

60T = 21T+3.5T2

3.5T =39

T=11.143 s